Question 9 (4 points)

Below are returns on the stock A and S&P500 index. All numbers are in decimals (-0.0222 is equivalent to -2.22%).

A S_P500
-0.0222 0.0032
-0.0048 -0.0058
0.1333 0.0434
0.0765 0.1081
-0.0161 -0.0121
0.1250 0.1400
0.0145 0.0368
-0.0475 -0.0454
0.0430 0.0577
-0.0260 -0.1374
0.0071 0.0064
0.0249 0.0186
0.0850 0.0215
-0.0624 -0.0752
0.0933 0.0365
0.0456 0.0528
-0.0632 -0.0131
0.0450 0.0009
0.0200 0.0017
0.0280 0.0985

Suppose that the next S&P500 return will be 7%. Use the CAPM model to calculate the expected return on the stock A. Assume the risk-free rate of 1%.

Please write an answer in decimals. For example, 12.34% would be 0.1234.
Also, round your answer to the fourth decimal.
Answer #1
A B C D E F G H
2
3 A S_P500
4 -2.22% 0.32%
5 -0.48% -0.58%
6 13.33% 4.34%
7 7.65% 10.81%
8 -1.61% -1.21%
9 12.50% 14.00%
10 1.45% 3.68%
11 -4.75% -4.54%
12 4.30% 5.77%
13 -2.60% -13.74%
14 0.71% 0.64%
15 2.49% 1.86%
16 8.50% 2.15%
17 -6.24% -7.52%
18 9.33% 3.65%
19 4.56% 5.28%
20 -6.32% -1.31%
21 4.50% 0.09%
22 2.00% 0.17%
23 2.80% 9.85%
24 Mean Return 2.50% 1.69% =AVERAGE(E4:E23)
25 St. Deviation 5.73% 6.23% =STDEV.S(E4:E23)
26 Cov(A,S&P 500) 0.002555
27
28 Beta for A =Cov(A,S&P 500) / Variance of Market
29 0.66 =D26/ (E25^2)
30
31 As Per CAPM, expected return can be calculated as
32 r(E) = rf + β*(rm-rf)
33 Using the Following data
34 Beta (β) 0.66
35 Risk free rate ( rf ) 1.0%
36 Market Return (rm) 7%
37
38 expected return can be calculated as follows:
39 Expected return = rf + β*(rm-rf)
40 =1%+0.66*(7%-1%)
41 4.95% =D35+D34*(D36-D35)
42
43 Hence expected return is 4.95%
44